# Missing Data Imputation by Reducing Mutual Information with Rectified Flows

### v2

@@ Abstract
This paper introduces a novel iterative method for missing data imputation that sequentially reduces the mutual information between data and their corresponding missing mask. Inspired by GAN-based approaches, which train generators to decrease the predictability of missingness patterns, our method explicitly targets the reduction of mutual information. Specifically, our algorithm iteratively minimizes the KL divergence between the joint distribution of the imputed data and missing mask, and the product of their marginals from the previous iteration. We show that the optimal imputation under this framework corresponds to solving an ODE, whose velocity field minimizes a rectified flow training objective. We further illustrate that some existing imputation techniques can be interpreted as approximate special cases of our mutual-information-reducing framework. Comprehensive experiments on synthetic and real-world datasets validate the efficacy of our proposed approach, demonstrating superior imputation performance. The code of the proposed algorithm can be found at https://github.com/anewgithubname/MIRI.
@@ 1 Introduction
The problem of missing data, referring to absent components (NaNs) in a dataset, can arise in a wide range of scenarios (Rubin, 1976; Little and Rubin, 2019). For example, in a survey dataset, respondents might skip some questions accidentally or intentionally, resulting in missing values. fMRI data could have missing voxels due to incomplete brain coverage and spatial variations in images acquired across subjects, (Vaden et al., 2012). Simple “one-shot” methods replace missing entries with summary statistics, such as the sample mean or median of the observed data. In contrast, approaches like Multiple Imputation by Chained Equations (MICE) (van Buuren and Groothuis-Oudshoorn, 2011), MissForest (Stekhoven and Bühlmann, 2012), and HyperImpute (Jarrett et al., 2022) employ an iterative, dimension-wise scheme. At each iteration, they model one feature conditional on all others and draw from that model to replace the feature’s missing entries. They cycle through the features repeatedly until the imputations stabilize. Because the updates are sequential and applied one dimension at a time, these methods are difficult to parallelize and scale poorly to high-dimensional data such as images (Brini et al., 2024). Their accuracy is also highly sensitive to the choice of per-variable models (van Buuren and Groothuis-Oudshoorn, 2011; Laqueur et al., 2022).
@@
In recent years, research on generative modeling has made huge progress. Since imputation is naturally a data-generating task, generative models have been applied to impute missing data in some recent works. (Uehara et al., 2020) proposes to maximize the likelihood of the imputed data using an EM algorithm, where the likelihood function is modelled using a normalizing flow (Rezende and Mohamed, 2015; Dinh et al., 2015). Generative Adversarial Imputation Nets (GAIN) (Yoon et al., 2018) leverages Generative Adversarial Nets (GAN) (Goodfellow et al., 2014) to impute missing components. It trains a generator (a neural network) so that a binary classifier cannot distinguish whether a given component is imputed or not. (Li et al., 2019) proposes MisGAN to train a generator so that a discriminator cannot differentiate between generated data and observed data through missing masks.
@@
Both GANs and normalizing flows come with their own challenges. GANs rely on unstable adversarial training and may suffer from mode collapse (Li et al., 2018; Zhang et al., 2018), while normalizing flows need specially designed architectures to ensure their invertibility, limiting their applications. Currently, diffusion models and flow-based methods represent the state of the art in generative modeling techniques (Song et al., 2021; Lipman et al., 2023; Liu et al., 2023). They first train a velocity field according to some criterion. Then, they obtain samples from the target distribution by solving an ODE or SDE using the trained velocity field as the drift. The main challenge is that the flow-based methods are designed to transport reference samples to match the target distribution. However, it is unclear how to formulate the missing data imputation problem as a distribution transport problem. Nonetheless, some progress along this line has been made. For example, MissDiff Ouyang et al. (2023) learns a conditional score model from partially observed samples, then uses an inpainter (Lugmayr et al., 2022) to impute missing values.
@@
Authors in (Liu et al., 2024) observe that GAIN’s criterion encourages the imputed dataset to be independent of the missingness pattern. In this paper, we continue their path to explore this idea but under a more rigorous framework. Our contributions are summarized as follows: (1) We introduce a novel imputation framework, named Mutual Information Reducing Iterations (MIRI), and prove that this iterative algorithm reduces the mutual information between imputed data and their corresponding missingness mask, when an optimal imputer is employed; (2) We show that an optimal MIRI imputer can be obtained by solving an ordinary differential equation (ODE), where the velocity field is trained using a rectified flow objective. This formulation naturally integrates flow-based generative modeling into the imputation process; (3) We show that several existing imputation methods can be viewed as approximate special cases of the MIRI framework; (4) We demonstrate that our proposed approach achieves promising empirical performance on both tabular and image datasets.
@@ 2 Background
We now briefly review the missing data imputation problem, the Generative Adversarial Imputation Nets (GAIN) (Yoon et al., 2018) and Rectified Flow (Liu et al., 2023), which are the foundations of our proposed method.
Notations: We denote scalars as lowercase letters (e.g., $x,y$), and vectors as bold lowercase letters (e.g., $\mathbf{x},\mathbf{y}$). Random variables are uppercase letters, e.g., $X,Y$. Random vectors are bold uppercase letters $\mathbf{X},\mathbf{Y}$. The superscript $t\in\{1,\dots,T\}$ denotes the iteration count in the algorithm. The subscript $0,1$ and $\tau$ denote the continuous time index in an ODE. Given a fixed missing mask $\mathbf{m}$ and a vector $\mathbf{x}$ with the same dimension, we write $\mathbf{x}_{\mathbf{m}}$ as the subvector of $\mathbf{x}$ whose elements correspond to $m_{i}=1$ and $\mathbf{x}_{1-\mathbf{m}}$ as the subvector of $\mathbf{x}$ whose elements corresponds to $m_{i}=0$. Following the convention, we refer to $\mathbf{x}_{\mathbf{m}}$ as the “non-missing components” of $\mathbf{x}$ and $\mathbf{x}_{1-\mathbf{m}}$ as the “missing components” of $\mathbf{x}$. $A\stackrel{{\scriptstyle d}}{{=}}B$ means random variables $A$ and $B$ are equal in distribution.
@@ 2.1 Sequential Imputation
Now, we formally define the missing data imputation problem. The true data vector $\mathbf{X}^{*}=[X^{*}_{1},\dots,X^{*}_{d}]$ is a random vector taking values on $\mathcal{X}\subset\mathbb{R}^{d}$, while the missing mask $\mathbf{M}=[M_{1},\dots,M_{d}]$ is a random vector in $\{0,1\}^{d}$. In the missing completely at random (MCAR) Rubin (1976) setting, we assume that the true data is independent of the missing mask, i.e., $\mathbf{X}^{*}\perp\!\!\!\!\perp\mathbf{M}$. The observed data vector $\tilde{\mathbf{X}}=(\tilde{X}_{1},\dots,\tilde{X}_{d})$ is defined as $\tilde{X}_{j}:=\begin{cases}X^{*}_{j}&M_{j}=1\\ \mathrm{NaN}&M_{j}=0\end{cases}.$ In words, the variable $j$ in the observation $\tilde{\mathbf{X}}$ is missing when $M_{j}=0$ and is observed when $M_{j}=1$. Equivalently,
$\tilde{\mathbf{X}}=\mathbf{M}\odot\mathbf{X}^{*}+(1-\mathbf{M})\odot\mathrm{NaN},$
@@
where $\mathbf{g}:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is called an imputer. $\mathbf{Z}_{0}$ is a random seed of the imputer. Different random seeds $\mathbf{Z}_{0}$ lead to different imputations.
In many cases, the imputation is a sequential process: Instead of computing $\mathbf{X}$ in one go, we improve the imputation by iterative updates. The sequential imputation considers the following procedure. For iteration $t=1,\dots,T$, we construct the imputation
@@
where $\mathbf{X}^{(t)}$ indicates the imputed data vector at iteration $t$. $\mathbf{X}^{(0)}$ is initialized with any standard imputation technique (e.g., mean or median imputation). We show a schematic of the sequential imputation algorithm in Figure 8 of Appendix G.
The central task of this iterative imputation is to learn the imputer $\mathbf{g}_{t}$ at each iteration $t$. In the following sections, we show how a popular missing data imputation scheme can be formulated within this framework and how $\mathbf{g}_{t}$ can be trained using a GAN-type objective.
2.2 Generative Adversarial Imputation Nets (GAIN) (Yoon et al., 2018)
GAIN arises from a natural intuition: With perfect imputation, imputed entries are indistinguishable from the originally observed ones. Thus, GAIN trains a generator so that a classifier cannot predict the missing mask $\mathbf{M}$ accurately.
Let $f(m_{j},\mathbf{x},\mathbf{m}_{-j})\in[0,1]$ be a probabilistic classifier which models the conditional probability $\mathbb{P}_{M_{j}\mid\mathbf{X}^{(t-1)},\mathbf{M}_{-j}}$, the probability of $j$-th component being imputed given the previous imputation $\mathbf{X}^{(t-1)}$ and the missing mask excluding the $j$-th component. For iteration $t=1,\dots,T$, GAIN performs the following steps:
@@
    $f_{tj}:=\arg\min_{f}-\mathbb{E}\left[\log f\left(M_{j},\mathbf{X}^{(t-1)},\mathbf{M}_{-j}\right)\right]$, for all $j=1\dots d$.
@@
    $\mathbf{g}_{t}:=\arg\max_{\mathbf{g}}-\sum_{j}\mathbb{E}\left[\log f_{tj}\left(M_{j},\mathbf{X}^{(t)}(\mathbf{g}),\mathbf{M}_{-j}\right)\right]-\lambda(\text{reconstruction error})$.
@@
$\mathbf{X}^{(t)}(\mathbf{g}_{t})$ is the imputer we introduced in (2). At iteration $t$, we first train classifiers $f_{tj}$ by minimizing the cross-entropy loss. Subsequently, we train the imputer $\mathbf{g}_{t}$, parameterized as a neural network, to reduce the classifiers’ predictive accuracy by maximizing the sum of cross-entropy losses across all $j$, while simultaneously minimizing the reconstruction error.
In practice, both the minimization and maximization steps are performed by a single gradient update, leading to an adversarial training. Thus, the algorithm is named Generative Adversarial Imputation Nets (GAIN). We observe that GAIN works well in practice. However, the adversarial training is known to be difficult and may suffer from mode collapses (Li et al., 2018; Zhang et al., 2018). Moreover, balancing the reconstruction error and the cross entropy loss requires careful tuning of the hyperparameter $\lambda$.
Recently, an observation was made that the generator training of GAIN may be viewed as a process of breaking the dependency between $\mathbf{M}$ and $\mathbf{X}^{(t)}(\mathbf{g})$ (Liu et al., 2024). Indeed, if $\mathbf{X}^{(t)}(\mathbf{g})\perp\!\!\!\!\perp\mathbf{M}$, implying $\mathbf{X}^{(t)}(\mathbf{g})$ is not predictive of $\mathbf{M}$ at all, then the cross-entropy loss is maximized. In this paper, we rigorously explore the idea of imputation by reducing dependency, which will guide us in developing a novel iterative imputation algorithm that reduces the dependency between $\mathbf{M}$ and $\mathbf{X}^{(t)}(\mathbf{g})$.
@@ 2.3 Flow-based Sampling and Rectified Flow
As an alternative to GANs, flow-based generative models have attracted significant interest in recent years. These algorithms first train a “velocity field” over time $\tau\in[0,1]$ according to some loss function, and then generate samples by solving the ODE or SDE using the velocity field as the drift. In some cases, the loss function is a simple least-squares loss, leading to a training routine that is simpler and more stable than GAN. Methods in this category include Score-based generative models (Song and Ermon, 2019; Song et al., 2021), Rectified Flow (Liu et al., 2023), Flow Matching (Lipman et al., 2023), Duffusion Schrödinger Bridge (De Bortoli et al., 2021) , among others.
Rectified flow is one of the simplest flow-based generative models. Given two random vectors $\mathbf{X}_{0}$ and $\mathbf{X}_{1}$, the velocity field $\mathbf{v}^{*}$ is trained by minimizing the following objective:
@@
where $\mathbf{X}_{\tau}$ is the interpolated random vector between $\mathbf{X}_{0}$ and $\mathbf{X}_{1}$, defined as $\mathbf{X}_{\tau}=\tau\mathbf{X}_{1}+(1-\tau)\mathbf{X}_{0}$ for $\tau\in[0,1]$. The sampling ODE is: $\frac{\mathrm{d}\mathbf{Z}_{\tau}}{\mathrm{d}\tau}=\mathbf{v}^{*}(\mathbf{Z}_{\tau},\tau).$ One can prove that, if $\mathbf{Z}_{0}\stackrel{{\scriptstyle d}}{{=}}\mathbf{X}_{0}$, then $\mathbf{Z}_{\tau}\stackrel{{\scriptstyle d}}{{=}}\mathbf{X}_{\tau}$ for all $\tau\in[0,1]$. This “marginal preserving property” implies: If one initializes the ODE using a sample $\mathbf{X}_{0}$ from the reference distribution, one can obtain a sample from the target distribution by solving the ODE at $\tau=1$.
3 Reducing Dependency
In this section, we propose the MIRI imputation framework, obtain the optimal imputer under this framework using rectified flow, and theoretically justify its validity.

### v3

@@ Abstract
This paper introduces a novel iterative method for missing data imputation that sequentially reduces the mutual information between data and the corresponding missingness mask. Inspired by GAN-based approaches that train generators to decrease the predictability of missingness patterns, our method explicitly targets this reduction in mutual information. Specifically, our algorithm iteratively minimizes the KL divergence between the joint distribution of the imputed data and missingness mask, and the product of their marginals from the previous iteration. We show that the optimal imputation under this framework can be achieved by solving an ODE whose velocity field minimizes a rectified flow training objective. We further illustrate that some existing imputation techniques can be interpreted as approximate special cases of our mutual-information-reducing framework. Comprehensive experiments on synthetic and real-world datasets validate the efficacy of our proposed approach, demonstrating its superior imputation performance. Our implementation is available at https://github.com/yujhml/MIRI-Imputation.
@@ 1 Introduction
The problem of missing data, referring to absent components (NaNs) in a dataset, can arise in a wide range of scenarios (rubin1976inference; little2019statistical). For example, in a survey dataset, respondents might skip some questions accidentally or intentionally, resulting in missing values. In fMRI data, there can be missing voxels due to incomplete brain coverage and spatial variations in images acquired across subjects (vaden2012multiple). Simple “one-shot” methods replace missing entries with summary statistics, such as the sample mean or median of the observed data. In contrast, approaches like MICE (vanbuuren2011mice), MissForest (stekhoven2012missforest), and HyperImpute (Jarrett2022HyperImpute) employ an iterative, dimension-wise scheme. In each iteration, they model one feature conditional on all others and sample from that model to replace the feature’s missing entries, cycling through all features repeatedly until the imputations stabilize. Because their updates are sequential and dimension-wise, these methods are difficult to parallelize and scale poorly to high‑dimensional data such as images (brini_missing_2024). Their accuracy is also highly sensitive to the choice of per‑variable models (vanbuuren2011mice; laqueur_supermice_2022).
@@
In recent years, research on generative modeling has made huge progress. Since imputation is naturally a data-generating task, generative models have been increasingly applied to impute missing data. (Richardson_2020) proposes to maximize the likelihood of the imputed data using an EM algorithm, where the likelihood function is modeled using a normalizing flow (rezende2015variational; dinh_nice_2015). GAIN (yoon2018gain) leverages Generative Adversarial Nets (GANs) (Goodfellow2014) to impute missing components. It trains a generator (a neural network) so that a binary classifier cannot distinguish whether a given component is imputed or not. However, both GANs and normalizing flows come with their own challenges. GANs rely on unstable adversarial training and may suffer from mode collapse (li_limitations_2018; zhang_convergence_2018), while normalizing flows require specially designed architectures to ensure their invertibility, limiting their applications.
@@
Currently, diffusion models and flow-based methods represent the state of the art in generative modeling techniques (song2021scorebased; lipman2023flow; liu2023flow). They first train a velocity field according to some criteria. Then, they sample from the target distribution by solving an ODE or SDE using the trained velocity field as the drift. The main challenge to adapt these methods for missing value imputation is that most flow-based methods are designed to transport reference samples to match the target distribution. However, it is unclear how to formulate the missing data imputation problem as a distribution transport problem. Nonetheless, progress has been made in this area. For example, MissDiff (ouyang_missdiff_2023) learns a score model from partially observed samples, then uses an inpainter (Lugmayr_2022_CVPR) to impute missing values.
Recently, (liu24minimizing) observes that GAIN’s criterion encourages the imputed dataset to be independent of the missingness pattern. In this paper, we further explore this idea within a more rigorous framework. Our contributions are summarized as follows:
@@
1. We introduce Mutual Information Reducing Iterations (MIRI), a novel framework that provably reduces mutual information between imputed data and their missingness mask when using an optimal imputer; 2. We show that an optimal MIRI imputer can be obtained by solving an ODE whose velocity field is trained by a rectified flow objective, naturally integrating generative modeling; 3. We reveal that several existing imputation methods can be viewed as approximate special cases of the MIRI framework; 4. We demonstrate that our proposed approach achieves promising empirical performance on both tabular and image datasets.
@@ 2 Background
We now briefly review the missing data imputation problem, Generative Adversarial Imputation Nets (GAIN) (yoon2018gain), and Rectified Flow (liu2023flow), which form the foundation of our proposed method.
Notations.  We denote scalars with lowercase letters (e.g., $x,y$) and vectors with bold lowercase letters (e.g., $\mathbf{x},\mathbf{y}$). Random variables are denoted by uppercase letters (e.g., $X,Y$), and random vectors by bold uppercase letters (e.g., $\mathbf{X},\mathbf{Y}$). The superscript $t\in\{1,\dots,T\}$ denotes the iteration count of the algorithm. The subscripts $0,1,$ and $\tau$ denote the continuous time index in an ODE. Given a fixed missingness mask $\mathbf{m}$ and a vector $\mathbf{x}$ of the same dimension, we write $\mathbf{x}_{\mathbf{m}}$ as the subvector of $\mathbf{x}$ whose elements correspond to entries where $m_{i}=1$, and $\mathbf{x}_{1-\mathbf{m}}$ as the subvector whose elements correspond to entries where $m_{i}=0$. Following convention, we refer to $\mathbf{x}_{\mathbf{m}}$ as the “non-missing components” of $\mathbf{x}$ and $\mathbf{x}_{1-\mathbf{m}}$ as the “missing components” of $\mathbf{x}$. The notation $A\mathrel{\smash{\overset{\text{d}}{=}}}B$ means that the random variables $A$ and $B$ are equal in distribution.
@@ 2.1 Sequential Imputation
Now, we formally define the missing data imputation problem. The true data vector $\mathbf{X}^{*}=[X^{*}_{1},\dots,X^{*}_{d}]$ is a random vector taking values on $\mathcal{X}\subset\mathbb{R}^{d}$, while the missingness mask $\mathbf{M}=[M_{1},\dots,M_{d}]$ is a random vector in $\{0,1\}^{d}$. In the Missing Completely at Random (MCAR) (rubin1976inference) setting, we assume that the true data is independent of the missingness mask, i.e., $\mathbf{X}^{*}\perp\!\!\!\!\perp\mathbf{M}$ (See Section F for discussions on Missing At Random (MAR) setting). The observed data vector $\tilde{\mathbf{X}}=(\tilde{X}_{1},\dots,\tilde{X}_{d})$ is defined as $\tilde{X}_{j}:=\begin{cases}X^{*}_{j}&M_{j}=1\\ \texttt{NaN}&M_{j}=0\end{cases}.$ In words, the variable $j$ in the observation $\tilde{\mathbf{X}}$ is missing when $M_{j}=0$ and is observed when $M_{j}=1$. Equivalently,
$\tilde{\mathbf{X}}=\mathbf{M}\odot\mathbf{X}^{*}+(1-\mathbf{M})\odot\texttt{NaN},$ (1)
@@
[…]
In many cases, imputation is an iterative process: Instead of computing $\mathbf{X}$ in a single step, the imputation is improved through iterative updates. This sequential imputation involves the following procedure: for each iteration $t=1,\dots,T$, we construct the imputation
@@
where $\mathbf{X}^{(t)}$ indicates the imputed data vector at iteration $t$. The initial vector, $\mathbf{X}^{(0)}$ is initialized with any standard imputation technique (e.g., Gaussian noise, mean or median imputation). A schematic of this sequential imputation algorithm is provided in Figure 4 of Appendix A.
The central task of this iterative process is to learn the imputer $\mathbf{g}_{t}$ at each iteration $t$. In the following sections, we show how a popular missing data imputation scheme can be formulated within this framework and how $\mathbf{g}_{t}$ can be trained using a GAN-type objective.
2.2 Generative Adversarial Imputation Nets (GAIN) (yoon2018gain)
GAIN arises from a natural intuition: if an imputation is perfect, the imputed entries should be indistinguishable from the originally observed ones. Thus, GAIN trains a generator so that a classifier cannot predict the missingness mask $\mathbf{M}$ accurately.
Let $f(m_{j},\mathbf{x},\mathbf{m}_{-j})\in[0,1]$ be a probabilistic classifier modeling the conditional probability $\mathbb{P}_{M_{j}|\mathbf{X}^{(t-1)},\mathbf{M}_{-j}}$, which is the probability that the $j$-th component is observed given the imputed data from the previous step, $\mathbf{X}^{(t-1)}$, and the rest of the missingness mask, $\mathbf{M}_{-j}$.
For each iteration $t=1,\dots,T$, GAIN performs the following steps:
@@ 2.2 Generative Adversarial Imputation Nets (GAIN) (yoon2018gain)
    Classifier $f_{tj}:=\arg\min_{f}-\mathbb{E}\left[\log f\left(M_{j},\mathbf{X}^{(t-1)},\mathbf{M}_{-j}\right)\right]$, for all $j=1\dots d$.
@@
    Imputer $\mathbf{g}_{t}:=\arg\max_{\mathbf{g}}-\sum_{j}\mathbb{E}\left[\log f_{tj}\left(M_{j},\mathbf{X}^{(t)}(\mathbf{g}),\mathbf{M}_{-j}\right)\right]-\lambda(\text{reconstruction error})$.
@@
    Update $\mathbf{X}^{(t)}\stackrel{{\scriptstyle\text{impute}}}{{\leftarrow}}\mathbf{X}^{(t)}(\mathbf{g}_{t}),\ t\leftarrow t+1$.
Here, $\mathbf{X}^{(t)}(\mathbf{g})$ is the imputed vector from the imputer $\mathbf{g}$ as defined in (3). At iteration $t$, we first train classifiers $f_{tj}$ by minimizing the cross-entropy loss. Subsequently, we train the imputer $\mathbf{g}_{t}$, parameterized as a neural network, to reduce the classifiers’ predictive accuracy by maximizing the sum of cross-entropy losses across all $j$, while simultaneously minimizing the reconstruction error.
In practice, the minimization and maximization steps are performed via alternating gradient updates, leading to adversarial training. We observe that GAIN works well in practice. However, the adversarial training is known to be difficult and may suffer from mode collapses (li_limitations_2018; zhang_convergence_2018). Moreover, balancing the reconstruction error and the cross entropy loss requires careful tuning of the hyperparameter $\lambda$.
Recently, it was observed that the generator training of GAIN can be viewed as a process of breaking the dependency between $\mathbf{M}$ and $\mathbf{X}^{(t)}(\mathbf{g})$ (liu24minimizing). Indeed, if $\mathbf{X}^{(t)}(\mathbf{g})\perp\!\!\!\!\perp\mathbf{M}$, meaning $\mathbf{X}^{(t)}(\mathbf{g})$ is not predictive of $\mathbf{M}$ at all, then the cross-entropy loss is maximized. In this paper, we rigorously explore this idea of imputation by dependency reduction. Guided by this principle, our algorithm explicitly aims to reduce the dependency between $\mathbf{M}$ and $\mathbf{X}^{(t)}(\mathbf{g})$.
@@ 2.3 Flow-based Sampling and Rectified Flow
As an alternative to GANs, flow-based generative models have attracted significant interest in recent years. These algorithms first train a “velocity field” over time $\tau\in[0,1]$ according to some loss function, and then generate samples by solving the ODE or SDE using the learned velocity field as the drift. In some cases, the loss function is a simple least-squares loss, leading to a training routine simpler and more stable than GAN. Methods in this category include Score-based generative models (song2019generative; song2021scorebased), Rectified Flow (liu2023flow), Flow Matching (lipman2023flow), Diffusion Schrödinger Bridge (de2021diffusion) , among others.
Rectified flow is one of the simplest flow-based generative models. Given two random vectors $\mathbf{X}_{0}$ (from a reference distribution) and $\mathbf{X}_{1}$ (from a target distribution), the velocity field $\mathbf{v}^{*}$ is trained by minimizing the following objective:
@@
where $\mathbf{X}_{\tau}$ is the linear interpolation between $\mathbf{X}_{0}$ and $\mathbf{X}_{1}$, defined as $\mathbf{X}_{\tau}=\tau\mathbf{X}_{1}+(1-\tau)\mathbf{X}_{0}$ for $\tau\in[0,1]$. Samples are then generated by solving the ODE:
$\frac{\mathrm{d}\mathbf{Z}_{\tau}}{\mathrm{d}\tau}=\mathbf{v}^{*}(\mathbf{Z}_{\tau},\tau).$ (5)
One can prove that, if $\mathbf{Z}_{0}\mathrel{\smash{\overset{\text{d}}{=}}}\mathbf{X}_{0}$, then $\mathbf{Z}_{\tau}\mathrel{\smash{\overset{\text{d}}{=}}}\mathbf{X}_{\tau}$ for all $\tau\in[0,1]$. This “marginal-preserving property” guarantees that solving the ODE from $\tau=0$ to $\tau=1$ transports samples from the reference distribution to the target distribution.
3 Reducing Dependency with Mutual Information
In this section, we introduce the MIRI imputation framework, obtain the optimal imputer under this framework using rectified flow, and theoretically justify its validity under the Missing Completely at Random (MCAR) setting. The validity for the Missing at Random (MAR) scenario is in Appendix F.
Algorithm 1 MIRI with Rectified Flow (Single Imputation)
[…]
2:  for $t=1$ to $T$ do
3:   Initialize a neural network model $\mathbf{v}$.
4:   for $n=1$ to NN do
[…]
9:   end for
[…]
11:  end for
12:  return $\{\mathbf{X}^{(T)}\}$